OpenStudy (anonymous):

2. You are driving home from school steadily at 95.0 km/h for 130 km. It then begins to rain and you slow to 65.0 km/h. You arrive home after driving 3.33 hours. (a) How far is your hometown from school? (b) What was your average speed?

OpenStudy (anonymous):

time taken to travel 130 km is t1= 130/95=1.37 h so t2=3.33-1.37= 1.96 h it's time take travel in rain. (a) so distance between hometown to school =95*t1+65*t2 or 130+65*t2 (b) average speed = total distance/ total time= (130+65*t2)/(t1+t2) i hope you can simplify it now.

OpenStudy (anonymous):

are you get it?